OpenStudy (blackops2_143):

What is the solution to the equation 2(M-8)=4(M+6) ? m =

pooja195 (pooja195):

Start by distributing 

OpenStudy (blackops2_143):

ook

OpenStudy (blackops2_143):

still need help

OpenStudy (danjs):

when you have a parenthesis of terms being multiplied by another in the front like those... this is the distribute pattern, multiply each term by the front and add them all up \[\large 2*(M-8) = 2*M + 2*(-8) = 2M - 16\] do that to both sides of the =